OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL What is the exact value of the expression 8 times square root of 15 + 5 + square root of 5 plus squareroot of 15

OpenStudy (anonymous):

\[8\sqrt{15}+5+\sqrt{5}+\sqrt{15}\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

okay, well the best way(in my opinion to do this) is to factor out the square roots

OpenStudy (anonymous):



OpenStudy (anonymous):

you can add them since they have the same \[\sqrt{5}\]but don;t change the \[\sqrt{5}\]just like \[\sqrt{5}+\sqrt{5}=2\sqrt{5}\]

OpenStudy (anonymous):

i still don't follow

OpenStudy (anonymous):

here: \[8\sqrt{5}+5+\sqrt{5}+\sqrt{5}\rightarrow 5+10\sqrt{5}\]\[8+1+1=10\]this has the same \[\sqrt{5}\]so just add the number

OpenStudy (anonymous):

so would it be \[8\sqrt{15} + 5\sqrt{15}\]

OpenStudy (anonymous):

@Lynncake

OpenStudy (anonymous):

\[8\sqrt{5}+5+2\sqrt{5}\]

OpenStudy (anonymous):

thats not one of my choices:(

OpenStudy (anonymous):

lol, sorry. i got 5 mixed with 15. wait. that should be \[9\sqrt{15}+5+\sqrt{5}\]

OpenStudy (anonymous):

oh haha thanks i understand it now!

OpenStudy (anonymous):

you're welcome @helpmeplease123451